SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 6,600 of these meals using 1,950 direct labor-hours. The company paid its direct labor workers a total of $27,300 for this work, or $14.00 per hour.

According to the standard cost card for this meal, it should require 0.30 direct labor-hours at a cost of $13.00 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 6,600 meals?

2. What is the standard labor cost allowed (SH × SR) to prepare 6,600 meals?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

(For requirements 3 and 4, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do no round intermediate calculations.)

1. Standard labor-hours allowed
2. Standard labor cost allowed
3. Labor spending variance
4. Labor rate variance
Labor efficiency variance

Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours.

In the most recent month, 200,000 items were shipped to customers using 8,900 direct labor-hours. The company incurred a total of $32,040 in variable overhead costs.

According to the company’s standards, 0.03 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.65 per direct labor-hour.

Required:

1. What is the standard labor-hours allowed (SH) to ship 200,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 200,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?

(For requirements 3 and 4, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do no round intermediate calculations.)

1. Standard quantity of labor-hours allowed
2. Standard variable overhead cost allowed
3. Variable overhead spending variance
4. Variable overhead rate variance
Variable overhead efficiency variance
Answer #1

1.

Standard hours allowed = 6600*0.30

= 1980

Standard cost allowed = 1980*13

= 25,740

Labor spending variance = 25,740 - 27,300

= 1560 U

Labor Rate Variance = (SR-AR) *AH

= (13 - 14)*1950

= 1950 U

Labor efficiency variance = (SH-AH) *SR

= (1980 - 1950)*13

= 390 Favourable